# Towards Context-Aware Domain Generalization: Representing Environments with Permutation-Invariant Networks

### 15 Dec 2023 v1

@@ C.1 Data Generation
The YouTube channel minutephysics also published a short descriptive video on this phenomenon in 2017 [38].
C.2 Training Details
@@
C.3 Non-Linear Models
In the experiments in Section 4.2, we considered linear models for our model and the baseline. In the following, we show results for the non-linear model class in Figure 7. We compare predictions of a baseline model and our model on all environments in Figure 8. We see that the extrapolation task fails in some cases as in environment 1. This is due to the mismatch of the considered model class and ground truth model.
@@ C.3 Non-Linear Models
Appendix D Experiment 2: Details
@@
Figure 9: Experiment 2. We generate four distinct domains synthetically. Notably, the background color within each domain follows a normal distribution. However, there are variations in the means across these domains Note that there is a huge overlap between the environments.
Data samples from different environments are depicted in Figure 9. The process of how inputs relate to outputs is described in Section 4.3.
@@
Appendix E Experiment 3: Details
To select between the baseline model and the invariant model, we are required to distinguish between ID and OOD data. Therefore, we follow the approach proposed in Section 2.7. We consider the $k$-nearest neighbors of the training set to compute the score $s_{\boldsymbol{\psi}}$ where $k=5$. Since we compare the scores elicited by features of the baseline model with the scores elicited by the features extracted by the set-encoder, we restricted both architectures to have the same feature dimension. To establish a threshold for distinguishing between ID and OOD samples, we designate samples with scores below the 95% quantile of the validation set as ID and those above as OOD (see Section 2.7 for details).
@@ Appendix E Experiment 3: Details
Appendix F Experiment 4 and 5: Details
For the BikeSharing dataset we consider a simple feed-forward neural network in all models. For the PACS as well as the OfficeHome dataset we consider features $g(\mathbf{X})$ that are kept fixed and not optimized. Here, we employ the Clip features proposed in [47]. The inference model, baseline model, and set-encoder are kept simple and employ only linear layers followed by ReLU activation functions. Given that Clip features considerably simplify the task, we performed a minimal hyper-parameter search and ensured that the inference model had a similar number of parameters as the baseline model. In all cases, the set-encoder is kept simple and its hyper-parameters are selected for optimal performance of the contextual environment predictor $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$.
@@ Appendix F Experiment 4 and 5: Details
Appendix G Comparison of Permutation-Invariant Architectures
@@
Therefore, we construct the toy dataset as described in Section C.1, but use the setting n_domains = 2 and vary the distance between environments for each experiment.
@@
Each architecture consists of a linear projection into a $64$-dimensional feature space, followed by a fully connected network with $3$ hidden layers, each containing $64$ neurons and a ReLU [1] activation. For the set-based methods, this is followed by the respective pooling. We choose $8$ heads for the attention-based model.
Finally, the output is linearly projected back into the $2$-dimensional logit space, where the loss is computed via cross-entropy [73].
@@
Figure 10: Comparison of different architectural choices for the permutation-invariant network in predicting the data’s originating environment. We consider various distances between environments and different set sizes $n$. As anticipated, the plots illustrate that smaller environment distances make it more challenging to differentiate between them. Moreover, with larger set size $n$, our ability to predict the environment label improves. Notably, the baseline model shows significantly poorer performance compared to the model utilizing contextual information in the form of a set input.
Appendix H Bike Sharing Dataset
This dataset, taken from the UCI machine learning repository[17], consists of over $17000$ hourly and daily counts of bike rentals between 2011 and 2012 within the Capital bikeshare system.
@@ Appendix H Bike Sharing Dataset
Similar to [49], we only consider the hourly rental data. We drop information about the concrete date and information about casual versus registered renters. We choose the season variable (spring, summer, fall, winter) as the environment and the bike rental count as the regression target. Since we deal with count data, we also apply square root transformation on the target similar to [49].

### 21 Feb 2024 v2

@@ E.1 Data Generation
The YouTube channel minutephysics also published a short descriptive video on this phenomenon in 2017 [minutephysics and Reich, 2017].
E.2 Training Details
@@
E.3 Non-Linear Models
In the experiments in subsection 4.2, we considered linear models for our model and the baseline. In the following, we show results for the non-linear model class in Figure 7. We compare predictions of a baseline model and our model on all environments in Figure 8. We see that the extrapolation task fails in some cases as in environment 1. This is due to the mismatch of the considered model class and ground truth model.
@@ E.3 Non-Linear Models
Appendix F Additional Experiment: Details
@@
Figure 9: Additional Experiment. We generate four distinct domains synthetically. Notably, the background color within each domain follows a normal distribution. However, there are variations in the means across these domains Note that there is a huge overlap between the environments.
Data samples from different environments are depicted in Figure 9. The process of how inputs relate to outputs is described in Appendix B.
@@
Appendix G Experiment 2: Details
To select between the baseline model and the invariant model, we are required to distinguish between ID and OOD data. Therefore, we follow the approach proposed in subsection 2.5. We consider the $k$-nearest neighbors of the training set to compute the score $s_{\boldsymbol{\psi}}$ where $k=5$. Since we compare the scores elicited by features of the baseline model with the scores elicited by the features extracted by the set-encoder, we restricted both architectures to have the same feature dimension. To establish a threshold for distinguishing between ID and OOD samples, we designate samples with scores below the 95% quantile of the validation set as ID and those above as OOD (see subsection 2.5 for details).
@@ Appendix G Experiment 2: Details
Appendix H Experiment 3 and 4: Details
For the BikeSharing dataset we consider a simple feed-forward neural network in all models. For the PACS as well as the OfficeHome dataset we consider features $g(\mathbf{X})$ that are kept fixed and not optimized. Here, we employ the Clip features proposed in Radford et al. [2021]. The inference model, baseline model, and set-encoder are kept simple and employ only linear layers followed by ReLU activation functions. Given that Clip features considerably simplify the task, we performed a minimal hyper-parameter search and ensured that the inference model had a similar number of parameters as the baseline model. In all cases, the set-encoder is kept simple and its hyper-parameters are selected for optimal performance of the contextual environment predictor $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$.
Table 5: Experiment 4. Performance comparison between our model and the baseline, broken down by target domain. We compare their performance in the ID and OOD setting (MSE), as well as their capability to detect a novel environment (AUROC). Both models fail in the OOD setting, but our model can detect with strong certainty when this is the case. We present the mean and standard deviation derived from 5 runs using different seeds for partitioning into training, validation, and test sets.
@@ Appendix H Experiment 3 and 4: Details
Appendix I Comparison of Permutation-Invariant Architectures
@@
Therefore, we construct the toy dataset as described in subsection E.1, but use the setting n_domains = 2 and vary the distance between environments for each experiment.
@@
Each architecture consists of a linear projection into a $64$-dimensional feature space, followed by a fully connected network with $3$ hidden layers, each containing $64$ neurons and a ReLU [Agarap, 2019] activation. For the set-based methods, this is followed by the respective pooling. We choose $8$ heads for the attention-based model.
Finally, the output is linearly projected back into the $2$-dimensional logit space, where the loss is computed via cross-entropy [Zhang and Sabuncu, 2018].
@@
Figure 10: Comparison of different architectural choices for the permutation-invariant network in predicting the data’s originating environment. We consider various distances between environments and different set sizes $n$. As anticipated, the plots illustrate that smaller environment distances make it more challenging to differentiate between them. Moreover, with a larger set size $n$, our ability to predict the environment label improves. Notably, the baseline model shows significantly poorer performance compared to the model utilizing contextual information in the form of a set input.
Appendix J Bike Sharing Dataset
This dataset, taken from the UCI machine learning repository [Fanaee-T, 2013], consists of over $17000$ hourly and daily counts of bike rentals between 2011 and 2012 within the Capital bikeshare system.
@@
Similar to Rothenhäusler et al. [2021], we only consider the hourly rental data. We drop information about the concrete date and information about casual versus registered renters. We choose the season variable (spring, summer, fall, winter) as the environment and the bike rental count as the regression target. Since we deal with count data, we also apply square root transformation on the target similar to Rothenhäusler et al. [2021].